With an increase in the thickness of the soft tissue by 4cm, the intensity of the x ray beam passing through the tissue decreased 9 times. Calculate the linear tissue absorption rate
1
Expert's answer
2018-12-26T05:53:11-0500

The intensity of the beam decreases according to the law: I=I_0 e^(-rl)

Here, I_0-initial beam intensity,r-linear tissue absortion rate,l-thickness


If we decrease the tissue by 4cm: I^'=I_0 e^(-r(l+4cm) )=I/9=(I_0 e^(-rl))/9

So, we get: e^(-4r)=1/9→r=ln9/4=0.55cm^(-1) - linear tissue absortion rate